# Detectability of Lensed Gravitational Waves in Matched-Filtering Searches

### v1

@@ V Methodology
We perform three injection campaigns using a time-domain matched-filtering search pipeline, GstLAL, to investigate how lensing impacts the detectability of gravitational-wave signals in searches. The distributions of the recovered source parameters, matched-filter SNR, the signal consistency test value, and the FAR inform the waveform deviation and noise fluctuations in matched-filtering searches, which explain changes in detectability when varying the source parameters. The details of the matched-filtering search setup and injection parameters are described in Tab. 1.
@@ V.1 Matched-filtering searches
In matched-filtering searches, accurate likelihood-ratio calculations for the injections require collecting sufficient noise background to evaluate the denominator in Eq. 9. Moreover, templates with different source parameters can recover the same injections due to noise fluctuations. These two conditions require a template bank with a parameter space sufficiently larger than the injections’ source parameters. We use representative 30-30 solar masses and spinless BBH as our source parameters, so the only key independent variables are the lensing parameters (strength of lensing) and source distances (strength of signals). This allows us to construct a smaller template bank to search for gravitational waves efficiently in our study following Privitera et al. (2014). The bank covers binary black hole masses with $m_{1}=[10M_{\odot},90M_{\odot}]$, $m_{2}=[10M_{\odot},40M_{\odot}]$ and include 14,606 templates, ensuring that the injection lives within the parameter space covered by this bank. Similar to GWTC-3, we use the SEOBNRv4ROM waveform model approximant Bohé et al. (2017) for our templates with pure quadrupole mode and aligned spin Abbott et al. (2024). The aligned component spins are distributed in the range $[-0.99,0.99]$. One potential caveat is the accuracy of the FAR computation when using a smaller template bank for the detectability study. As demonstrated in the results of the first injection campaign (Sec. VI.1), the FAR computation in our detectability study is primarily affected by the accuracy of the recovered source parameters, which means that a larger template bank simply permits greater scope for inaccurate recovery of source parameters at a higher computational cost. This justifies our use of a smaller template bank.
@@
We analyzed LIGO Hanford and Livingston data from the third observing run (O3) of the LVK collaboration. Our data includes real noises and spans from the first week of O3 with GPS time $1238166011$ s to $1238787954$ s. The data quality and glitch mitigation follows that of GWTC-3  Abbott et al. (2024). To assess the impact of lensing on matched-filter searches, we analyzed the distributions of recovered chirp mass, effective spin, matched-filter SNR, the autocorrelation-based signal-consistency test value $\xi^{2}$ in GstLAL, and the resultant FAR. The detection threshold for our searches is set as the representative $\rm{FAR}_{\rm{thr}}=1$ in $30$ days.
@@
Figure 5: The role of signal strength in the accuracy of the recovered source parameters of lensed gravitational waves. The panels show the histogram of the recovered source parameter at different source distances in our second injection campaign (see Tab. 1). Left: chirp mass $\mathcal{M}_{c}$. Right: effective spin $\chi_{\rm{eff}}$. The black dotted line represents the injected value of the $30-30$ $M_{\odot}$ spinless binary black hole signals. The unlensed injections are at 125 Mpc (blue). All lensed injection use impact parameter $y=0.01$ and redshifted lens mass $M_{\rm{Lz}}=10^{5}M_{\odot}$ at difference source distances: 125 Mpc (yellow), 625 Mpc (purple), 1250 Mpc (black). The recovered source parameter is more accurate at larger source distances (lower SNR) when the lensing distortion is significant.
@@ V.2 Injection Campaigns
We first conduct two injection campaigns to study how lensing and signal strength influence detectability. These allow us to predict the general picture of how lensing can impact the detectability of gravitational waves in matched-filtering searches in the third injection campaign. We prepare $11$ injection sets, each consisting of $19432$ injections. All injections are generated using the SEOBNRv4 Bohé et al. (2017) waveform approximant in LALSimulation LIGO Scientific Collaboration et al. (2018). We uniformly sample over the right ascension, declination, inclination, and polarization for all injections. All injections are distributed uniformly across the data over time. We summarize the lens parameters, the source distances, and the number of found injections in Tab. 1. In the following sections, we will discuss how we systematically vary the lens parameters and source distances for different injection sets to compare the role of lensing and signal strength on the detectability.
@@ V.2.1 Injection Campaign 1: The Role of Lensing
The first injection campaign explores the impact of lensing amplification on gravitational wave signal detectability. Four injection sets are examined: one unlensed set and three lensed sets with varying lensing parameters ($y=0.01$ and $M_{\rm{Lz}}=\{10,10^{3},10^{5}\}M_{\odot}$). These sets represent different lensing regimes: negligible amplification (long-wavelength), significant amplification without beating patterns (intermediate), and wave-dominated with notable beating patterns, as shown in Fig. 2. As lensing strength increases across sets, the match between lensed and unlensed waveforms decreases (=$\mathcal{M}=\{0.99,0.98,0.7\}$). As shown in Fig 1, conventional optimal SNR analysis suggests stronger gravitational lensing, which increases optimal SNR and should enhance signal detectability, which will be examined in this injection campaign.
@@ V.2.2 Injection Campaign 2: The Role of Signal Strength
The first injection campaign utilized a fixed unlensed SNR of $\rho\sim 100$, representing an unrealistically high SNR regime with O3 detector sensitivity and minimal impact from background noise. To better understand how signal strength influences lensing’s impact on detectability, a second injection campaign was conducted. This campaign used the same lens parameters as set 4, the one giving the largest lensing distortions ($y=0.01$ and $M_{\rm{Lz}}=10^{5}M_{\odot}$), and compared it to the unlensed case by varying source distances from $d_{L}=\{125,625,1250\}$ Mpc. This resulted in injections with unlensed optimal SNRs of $\rho_{\rm{opt}}\sim\{100,50,10\}$ at O3 sensitivity, corresponding to injection sets 1 and 4-8 in Tab. 1. Additionally, this investigation hints at whether improved detector sensitivities, leading to higher detector SNRs, could enhance the detectability of lensed gravitational waves.
Figure 6: The role of signal strength in the detectability of lensed gravitational waves. The panels show the distribution of the simulated events in our second injection campaign (see Tab. 1) as a function of matched-filter SNR $\rho$ and the signal consistency test $\xi^{2}/\rho^{2}$ (see Eq. 7). The color bar represents the false-alarm-rate (FAR). An event is considered a significant gravitational wave candidate when its FAR is below $3.85\times 10^{-7}$ Hz or 1 every 30 days, indicated by the red tick. The circle and triangular markers represent larger and smaller injections than the FAR threshold for the detection, respectively. Upper: Unlensed injections. Bottom: Lensed injections with impact parameter $y=0.01$,and redshifted lens mass $M_{Lz}=10^{5}M_{\odot}$. Left: $125$ Mpc, Middle: $625$ Mpc, Right: $1250$ Mpc. Gravitational waves with strong lensing distortion are less detectable at lower source distances (higher SNR).
@@ V.2.3 Injection Campaign 3: Detectability of Lensed Gravitational Waves
The initial two injection campaigns elucidated the roles of lensing and signal strength, enabling us to predict lensed gravitational wave detectability more comprehensively. To present a general picture of how lens parameters influence matched-filtering searches, we sampled over $y\in[0.01,1]$ and $M_{\rm{Lz}}\in[10M_{\odot},10^{5}M_{\odot}]$ for each injection set. We further verified the optimal SNR analysis by comparing matched-filter SNRs of lensed and unlensed injections, using sets with $d_{L}={125,625,1250}$ Mpc (corresponding to sets 1, 5, 6, and 9-11 in Tab. 1). These six injection sets also provide insights into the potential need for extending current template banks to include lensed gravitational waves by comparing their detectability in matched-filtering searches across different SNRs.
The first two injection campaigns (sets 1-8) used fixed component masses, spins, lens parameters, and source distances for each set. Distributing these injections across different times in the data allowed for systematic comparison of recovered $\mathcal{M}_{c}$ and $\chi_{\rm{eff}}$ under various noise realizations, providing specific insights into how lensing and signal strength impact matched-filter SNR. The third campaign, sampling over lens parameters, sacrificed these specific details for a more general picture of lensing’s impact on detectability, as each injection had different noise realizations and lens parameters. Consequently, recovered $\mathcal{M}c$ and $\chi_{\rm{eff}}$ are presented only for the first two campaigns, while matched-filter SNR, $\xi^{2}$, and FAR are reported for all campaigns.
Figure 7: Detectability of lensed gravitational waves. The panels show the distribution of the simulated events in our third injection campaign (see Tab. 1) as a function of the impact parameter $y$ and the redshifted lens mass $M_{\rm{Lz}}$. From top to bottom, the color bar represents the ratio of the lensed signal-to-noise ratio (SNR) $\rho_{\rm L}$ and unlensed SNR $\rho_{\rm UL}$, the signal consistency test $\xi^{2}/\rho^{2}$ (see Eq. 7), and the false-alarm-rate (FAR), respectively. An event is considered a significant gravitational wave candidate when its FAR is below $3.85\times 10^{-7}$ Hz or 1 every 30 days, indicated by the red tick. The source distances from left to right panels are $125$, $625$, and $1250$ Mpc. A large fraction of the lens parameter space does not pass the FAR threshold even at high SNR.
@@ VI Results
Our results highlight how lensing impacts the matched-filter SNR and the $\xi^{2}$ value. This influence scales with lensing amplification and signal strength, which is not predicted in optimal SNR analyses. In the following sections, we present such effects and show a general picture of the detectability of lensed gravitational waves.
@@ VI.1 The Role of Lensing
We compare the recovered distributions of $\mathcal{M}_{c}$ and $\chi_{\rm{eff}}$ for sets 1-4 in Fig. 3. At 125 Mpc, unlensed injections have an expected SNR of  100 at O3 sensitivity, with minimal noise impacting recovered template parameters. The unlensed (blue) and set 2 (red) injections show negligible differences, both centered at $\mathcal{M}_{c}=26.7M_{\odot}$ and $\chi_{\rm{eff}}=0$. For $M_{\rm{}_{Lz}}=10^{3}M_{\odot}$ (green), $\chi_{\rm{eff}}$ shifts to  0.25, while $\mathcal{M}_{c}$ remains qualitatively similar to unlensed injections. Set 4 (purple) shows significant lensing amplification, resulting in a substantial drop in recovered $\mathcal{M}_{c}$ and a multimodal $\chi_{\rm{eff}}$ distribution peaking at ${-1,0,1}$. This demonstrates that at $\rho\sim 100$, strong lensing distortion can significantly deviate recovered parameters from injected values, potentially reduces matched-filter SNR. However, the right panel in Fig. 1 expected an increase in SNR due to amplitide enhancement. Therefore, the resultant matched-filter SNR is a competition between the inaccuracy of recovered source parameters (reduce SNR) and lensing amplification (increase SNR).
@@
Fig. 4 illustrates the competing effects of lensing on detectability by showing $\xi^{2}/\rho^{2}$ vs $\rho$ for sets 1-4, with FAR values as the colormap. The unlensed scenario (upper left) demonstrates reduced FAR with higher SNR and decreasing $\xi^{2}/\rho^{2}$, with 17020 out of 19432 injections passing the detection threshold of $\rm{FAR}_{\rm{thr}}=3.85\times 10^{-7}$ Hz. Note that even in the representiaive unlensed $30-30$ $M_{\odot}$ spinless BBH at fixed source distance, sample over other source parameters and different noise realization can result in a wide rage of SNR value, subsequently lead to a wide range of $\rho$ from $10$ to $100$. Therefore, this illustrative unlensed injection set already show that the actual matched-filter SNR differ from optimal SNR. Moreove, small lensing distortion in the upper right panel shows a similar distribution with slightly more found injections (17068), attributed to minor signal amplification and negligible waveform distortion.
@@
The intermediate lensing distortion in the bottom left panel illustrates, for the first time, the competition between the inaccuracy of the recovered source parameters and lensing amplification. Although we observe an overall increase in SNR distribution due to lensing amplifcation, we also find a noticable increase in density of low SNR injections with $\rho<10$. Thus, the resultant SNR due to the above-mentioned competition depends on other factors, such as noise realizations and other source parameters being sampled in the injection set. After the SNR computation, the unlensed template that retrieved the lensed injections construct an expected SNR time series, while the actual SNR time series contain lensing distortion effect. Thus, calculating the $\xi^{2}$ in Eq. 7 will reflect this lensing distortion as an elevated $\xi^{2}$ value. Although $\xi^{2}$ is designed for mitigating loud noise triggers, it does not distinguish whether the deviation SNR time series is from the noises or the waveform distortion from the missing physics. Consequently, we can observe an overall increase in $\xi^{2}/\rho^{2}$, where the injections with higher SNR having the most significant enhancement from $\xi^{2}/\rho^{2}\sim 10^{-4}$ to $\xi^{2}/\rho^{2}\sim 10^{-3}$. Since $\rho$ and $\xi^{2}$ play a main role in FAR computation, we observe an overall enhancement in FAR, subsequenly reduce the number of found injections (14251).
@@
Finally, our injection set with the strongest lensing distortion presents an scenario where only $1$ out of $19421$ injections is found. First, a remarkable fraction of lensed injections is observed at $\rho_{\rm{opt}}<10$ due to poor inacurracy of recovered source parameters. This inaccurately recovered template further constructs an SNR time series for $\xi^{2}$ computation, thereby resulting in a $\xi^{2}$ value not intended for handling such large waveform distortion scenarios. This leads to a significant FAR increase from $<10^{-10}$ Hz to $>1$ Hz, resulting in only one found injection.
@@
These results demonstrate that significant waveform distortion impacts matched-filter SNR behavior, suggesting that optimal SNR poorly represents detectability even at $\rho_{\rm{opt}}\sim 100$. Furthermore, employing a large template bank with millions of templates in a realistic matched-filtering setup spans a broader source parameter space, but only yields a wider recovered source parameter space, leading to a more inaccurate SNR time series for $\xi^{2}$. Thus, our initial injection campaign results also support the use of a smaller template bank in detectability studies.

### v2

@@ V Methodology
We perform three injection campaigns using a time-domain matched-filtering search pipeline, GstLAL, to investigate how lensing impacts the detectability of gravitational-wave signals in searches. The distributions of the recovered source parameters, matched-filter SNR, the signal-consistency test value, and the FAR inform the waveform deviation and noise fluctuations in matched-filtering searches, which explain changes in detectability when varying the source parameters. The details of the matched-filtering search setup and injection parameters are described in Table 1.
@@ V.1 Matched-filtering searches
In matched-filtering searches, accurate likelihood-ratio calculations for the injections require collecting sufficient noise background to evaluate the denominator in Eq. (9). Moreover, templates with different source parameters can recover the same injections due to noise fluctuations. These two conditions require a template bank with a parameter space sufficiently larger than the injections’ source parameters. We use representative $30$-$30$ solar masses and a spinless binary black hole (BBH) as our source parameters, so the only key independent variables are the lensing parameters (strength of lensing) and source distances (strength of signals). This allows us to construct a smaller template bank to search for gravitational waves efficiently in our study following Privitera et al. (2014). The bank covers BBH masses with $m_{1}=[10M_{\odot},90M_{\odot}]$, $m_{2}=[10M_{\odot},40M_{\odot}]$ and includes 14606 templates, ensuring that the injection lives within the parameter space covered by this bank. Similar to the third Gravitational-Wave Transient Catalog GWTC-3, we use
seobnrv4_rom waveform model approximant Bohé et al. (2017) for our templates with pure quadrupole mode and aligned spin Abbott et al. (2024). The aligned component spins are distributed in the range $[-0.99,0.99]$. One potential caveat is the accuracy of the FAR computation when using a smaller template bank for the detectability study. As demonstrated in the results of the first injection campaign (Sec. VI.1), the FAR computation in our detectability study is primarily affected by the accuracy of the recovered source parameters, which means that a larger template bank simply permits greater scope for inaccurate recovery of source parameters at a higher computational cost. This justifies our use of a smaller template bank.
@@
We analyzed LIGO Hanford and Livingston data from the third observing run (O3) of the LVK Collaboration. Our data include real noises and spans from the first week of O3 with GPS time $1238166011$-$1238787954$ s. The data quality and glitch mitigation follows that of GWTC-3  Abbott et al. (2024). To assess the impact of lensing on matched-filter searches, we analyzed the distributions of recovered chirp mass, effective spin, matched-filter SNR, the autocorrelation-based signal-consistency test value $\xi^{2}$ in GstLAL, and the resultant FAR. The detection threshold for our searches is set as the representative $\rm{FAR}_{\rm{thr}}=1$ in $30$ days.
@@
Figure 5: The role of signal strength in the accuracy of the recovered source parameters of lensed gravitational waves. The panels show the histogram of the recovered source parameter at different source distances in our second injection campaign (see Table 1). Left: chirp mass $\mathcal{M}_{c}$. Right: effective spin $\chi_{\rm{eff}}$. The red solid line on the x-axis represents the injected value of the $30-30$ $M_{\odot}$ spinless binary black hole signals. The unlensed injections are at 125 Mpc (blue). All lensed injections use impact parameter $y=0.01$ and redshifted lens mass $M_{\rm{Lz}}=10^{5}M_{\odot}$ at difference source distances: 125 (purple), 625 (orange), 1250 Mpc (green). The recovered source parameter is more accurate at larger source distances (lower SNR) when the lensing distortion is significant.
@@ V.2 Injection campaigns
We first conduct two injection campaigns to study how lensing and signal strength influence detectability. These allow us to predict the general picture of how lensing can impact the detectability of gravitational waves in matched-filtering searches in the third injection campaign. We prepare $11$ injection sets, each consisting of $19432$ injections. All injections are generated using the
seobnrv4 Bohé et al. (2017) waveform approximant in LALSimulation LIGO Scientific Collaboration et al. (2018). We uniformly sample over the right ascension, declination, inclination, and polarization for all injections. All injections are distributed uniformly across the data over time. We summarize the lens parameters, the source distances, and the number of found injections in Table 1. In the following sections, we will discuss how we systematically vary the lens parameters and source distances for different injection sets to compare the role of lensing and signal strength on the detectability.
@@ V.2.1 Injection campaign 1: The role of lensing
The first injection campaign explores the impact of lensing amplification on gravitational-wave signal detectability. Four injection sets are examined: one unlensed set and three lensed sets with varying lensing parameters ($y=0.01$ and $M_{\rm{Lz}}=\{10,10^{3},10^{5}\}M_{\odot}$). These sets represent different lensing regimes: negligible amplification (long wavelength), significant amplification without beating patterns (intermediate), and wave dominated with notable beating patterns, as shown in Fig. 2. As lensing strength increases across sets, the match between lensed and unlensed waveforms decreases (=$\mathcal{M}=\{0.99,0.98,0.7\}$). As shown in Fig 1, conventional optimal SNR analysis suggests stronger gravitational lensing, which increases optimal SNR and should enhance signal detectability, which will be examined in this injection campaign.
@@ V.2.2 Injection campaign 2: The role of signal strength
The first injection campaign utilized a fixed unlensed SNR of $\rho\sim 100$, representing an unrealistically high-SNR regime with O3 detector sensitivity and minimal impact from background noise. To better understand how signal strength influences lensing’s impact on detectability, a second injection campaign was conducted. This campaign used the same lens parameters as set 4, the one giving the largest lensing distortions ($y=0.01$ and $M_{\rm{Lz}}=10^{5}M_{\odot}$), and compared it to the unlensed case by varying source distances from $d_{L}=\{125,625,1250\}$ Mpc. This resulted in injections with unlensed optimal SNRs of $\rho_{\rm{opt}}\sim\{100,50,10\}$ at O3 sensitivity, corresponding to injection sets 1 and $4$-$8$ in Table 1. Additionally, this investigation hints at whether improved detector sensitivities, leading to higher detector SNRs, could enhance the detectability of lensed gravitational waves.
Figure 6: The role of signal strength in the detectability of lensed gravitational waves. The panels show the distribution of the simulated events in our second injection campaign (see Table 1) as a function of matched-filter SNR $\rho$ and the signal-consistency test $\xi^{2}/\rho^{2}$ (see Eq. 7). The color bar represents the FAR. An event is considered a significant gravitational-wave candidate when its FAR is below $3.85\times 10^{-7}$ Hz or one every 30 days, indicated by the red tick. The circle and triangular markers represent larger and smaller injections than the FAR threshold for the detection, respectively. Upper: unlensed injections. Bottom: lensed injections with impact parameter $y=0.01$,and redshifted lens mass $M_{\rm Lz}=10^{5}M_{\odot}$. Left: $125$ Mpc. Middle: $625$ Mpc. Right: $1250$ Mpc. Gravitational waves with strong lensing distortion are less detectable at lower source distances (higher SNR).
@@ V.2.3 Injection campaign 3: detectability of lensed gravitational waves
The initial two injection campaigns elucidated the roles of lensing and signal strength, enabling us to predict lensed gravitational-wave detectability more comprehensively. To present a general picture of how lens parameters influence matched-filtering searches, we sampled over $y\in[0.01,1]$ and $M_{\rm{Lz}}\in[10M_{\odot},10^{5}M_{\odot}]$ for each injection set. We further verified the optimal SNR analysis by comparing matched-filter SNRs of lensed and unlensed injections, using sets with $d_{L}={125,625,1250}$ Mpc (corresponding to sets 1, 5, 6, and $9$-$11$ in Table 1). These six injection sets also provide insights into the potential need for extending current template banks to include lensed gravitational waves by comparing their detectability in matched-filtering searches across different SNRs.
The first two injection campaigns (sets $1$-$8$) used fixed component masses, spins, lens parameters, and source distances for each set. Distributing these injections across different times in the data allowed for systematic comparison of recovered $\mathcal{M}_{c}$ and $\chi_{\rm{eff}}$ under various noise realizations, providing specific insights into how lensing and signal strength impact matched-filter SNR. The third campaign, sampling over lens parameters, sacrificed these specific details for a more general picture of lensing’s impact on detectability, as each injection had different noise realizations and lens parameters. Consequently, recovered $\mathcal{M}_{c}$ and $\chi_{\rm{eff}}$ are presented only for the first two campaigns, while matched-filter SNR, $\xi^{2}$, and FAR are reported for all campaigns.
Figure 7: Detectability of lensed gravitational waves. The panels show the distribution of the simulated events in our third injection campaign (see Table 1) as a function of the impact parameter $y$ and the redshifted lens mass $M_{\rm{Lz}}$. From top to bottom, the color bar represents the ratio of the lensed SNR $\rho_{\rm L}$ and unlensed SNR $\rho_{\rm UL}$, the signal-consistency test $\xi^{2}/\rho^{2}$ (see Eq. 7), and the FAR, respectively. An event is considered a significant gravitational-wave candidate when its FAR is below $3.85\times 10^{-7}$ Hz or one every 30 days, indicated by the red tick. The source distances from left to right panels are $125$, $625$, and $1250$ Mpc. A large fraction of the lens parameter space does not pass the FAR threshold even at high SNR.
@@ VI Results
Our results highlight the impact of lensing on the matched-filter SNR and the $\xi^{2}$ value. This influence scales with lensing amplification and signal strength, an effect not accounted for in optimal SNR analyses. In the following sections, we explore these effects and provide a general overview of the detectability of lensed gravitational waves.
@@ VI.1 The role of lensing
We compare the recovered distributions of $\mathcal{M}_{c}$ and $\chi_{\rm{eff}}$ for sets $1$-$4$ in Fig.3. At 125 Mpc, unlensed injections have an expected SNR of approximately 100 at O3 sensitivity, with minimal noise affecting the recovered template parameters. The unlensed (blue) and set 2 (red) injections exhibit negligible differences, both centered at $\mathcal{M}_{c}=26.7M{\odot}$ and $\chi_{\rm{eff}}=0$. For $M_{\rm{Lz}}=10^{3}M_{\odot}$ (green), $\chi_{\rm{eff}}$ shifts to approximately 0.25, while $\mathcal{M}_{c}$ remains qualitatively similar to unlensed injections. Set 4 (purple) shows significant lensing amplification, resulting in a substantial drop in recovered $\mathcal{M}c$ and a multimodal $\chi{\rm{eff}}$ distribution peaking at ${-1,0,1}$. This demonstrates that at $\rho\sim 100$, strong lensing distortion can significantly deviate recovered parameters from injected values, potentially reducing the matched-filter SNR. However, the right panel in Fig.1 indicates an expected increase in SNR due to amplitude enhancement. Therefore, the resultant matched-filter SNR reflects a competition between the inaccuracy of recovered source parameters (reducing SNR) and lensing amplification (increasing SNR).
@@
Figure 4 illustrates the competing effects of lensing on detectability by showing $\xi^{2}/\rho^{2}$ vs $\rho$ for sets $1$-$4$, with FAR values represented as the color map. The unlensed scenario (upper left) demonstrates a reduced FAR with higher SNR and decreasing $\xi^{2}/\rho^{2}$, with 17020 out of 19432 injections passing the detection threshold of $\rm{FAR}_{\rm{thr}}=3.85\times 10^{-7}$ Hz. Note that even in the representative unlensed $3030$ $M_{\odot}$ spinless BBH at a fixed source distance, sampling over other source parameters and different noise realizations can result in a wide range of SNR values, subsequently leading to a wide range of $\rho$ values from $10$ to $100$. Thus, this illustrative unlensed injection set already demonstrates that the actual matched-filter SNR differs from the optimal SNR. Moreover, small lensing distortions in the upper right panel show a similar distribution with slightly more successful detections (17068), attributed to minor signal amplification and negligible waveform distortion.
@@
The intermediate lensing distortion in the bottom left panel illustrates, for the first time, the competition between the inaccuracy of the recovered source parameters and lensing amplification. Although we observe an overall increase in the SNR distribution due to lensing amplification, we also find a noticeable increase in the density of low-SNR injections with $\rho<10$. Thus, the resultant SNR, influenced by the above-mentioned competition, depends on other factors such as noise realizations and the source parameters sampled in the injection set. After the SNR computation, the unlensed template used to retrieve the lensed injections constructs an expected SNR time series, while the actual SNR time series contains lensing distortion effects. As a result, calculating $\xi^{2}$ using Eq.( 7) reflects this lensing distortion as an elevated $\xi^{2}$ value. Although $\xi^{2}$ is designed to mitigate loud noise triggers, it does not distinguish whether deviations in the SNR time series arise from noise or waveform distortions caused by missing physics. Consequently, we observe an overall increase in $\xi^{2}/\rho^{2}$, with the injections at higher SNR exhibiting the most significant enhancement, increasing from $\xi^{2}/\rho^{2}\sim 10^{-4}$ to $\xi^{2}/\rho^{2}\sim 10^{-3}$. Since $\rho$ and $\xi^{2}$ play a key role in FAR computation, this increase leads to an overall enhancement in FAR, subsequently reducing the number of successful detections to 14251.
@@
Our injection set with the strongest lensing distortion presents a scenario where only $1$ out of $19421$ injections is successfully detected. A remarkable fraction of lensed injections is observed at $\rho_{\rm{opt}}<10$ due to the poor accuracy of the recovered source parameters. This inaccurately recovered template further constructs an SNR time series for $\xi^{2}$ computation, resulting in a $\xi^{2}$ value that is not designed to handle such large waveform distortion scenarios. As a consequence, this leads to a significant increase in FAR, rising from $<10^{-10}$ to $>1$ Hz, and ultimately results in only one successful detection.
@@
These results demonstrate that significant waveform distortion impacts the behavior of matched-filter SNR, suggesting that optimal SNR poorly represents detectability even at $\rho_{\rm{opt}}\sim 100$. Furthermore, employing a large template bank with millions of templates in a realistic matched-filtering setup spans a broader source parameter space but only yields a wider recovered source parameter space. This broader recovered space leads to a more inaccurate SNR time series for $\xi^{2}$. Thus, our initial injection campaign results also support the use of a smaller template bank in detectability studies.
Fully understanding the inaccuracies in the recovered source parameters for the lensed injections, it is essential to conduct a detailed investigation of each template’s response. This includes analyzing metrics such as the SNR time series and the $\xi^{2}$ time series in response to the lensed injections. However, such an investigation is computationally intensive, requiring significant computational resources and time to process and analyze the data comprehensively. Consequently, this level of detail falls beyond the scope of the current study and is deferred to future work.